# Admissibility Breakdown in High-Dimensional Sparse Regression with $\ell_{1}$ Regularization

### v1

@@ 3 Main Results
Remark 3.10.
The preceding lemmas indicate that the structure of the design matrix affects the choice of the Lasso tuning parameter. In particular, when the columns of the design matrix are more strongly correlated, a smaller tuning parameter is required, since overly large tuning values may lead to inadmissibility. In addition, a direct calculation of the prediction improvement is possible; see Lemma A.5 for further details. Finally, we consider a special case in which the Lasso selects the true support with high probability, taking into account the symmetry of the noise in addition to its tail behavior.
@@
Let $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$. Consider our Lasso-Ridge method. Select the tuning parameter of the second step as $\lambda_{\mathrm{R}}=c|E|$, then we have
@@
This theorem shows that when the Lasso selects the true support with high probability, its prediction performance can be improved by a suitable refinement over a wide range of tuning parameters. Moreover, it is known that variable selection consistency may occur even when the Lasso tuning parameter does not vanish; see, for example, (Lahiri, 2021). In such settings, the present theorem implies that prediction improvement can be achieved.
4 Extensions
The analysis developed in this paper is not limited to the classical Lasso estimator. Its core arguments extend naturally to other $\ell_{1}$-based procedures, most notably the Dantzig selector and the Adaptive Lasso. We describe these extensions below.
Dantzig selector.
The Dantzig selector (Candès and Tao, 2007) is defined by
$\hat{\beta}_{\mathrm{D}}\in\operatorname*{argmin}_{\beta\in\mathbb{R}^{p}}\|\beta\|_{1}\quad\text{subject to}\quad\left\|X^{\top}(y-X\beta)/n\right\|_{\infty}\leq\lambda_{\mathrm{D}}.$
With the equalcorrelation set defined analogously under the Dantzig tuning parameter $\lambda_{\mathrm{D}}$, all arguments and results developed above continue to hold with only notational modifications.
Adaptive Lasso.
The Adaptive Lasso (Zou, 2006) solves
$\hat{\beta}_{\mathrm{A}}\in\operatorname*{argmin}_{\beta\in\mathbb{R}^{p}}\Bigl\{\|y-X\beta\|_{n}^{2}+\lambda_{\mathrm{A}}\sum_{j=1}^{p}w_{j}|\beta_{j}|\Bigr\},$
where $w_{j}>0$ are data-dependent weights, typically derived from an initial estimator. For columns of $X$ corresponding to a single tuning weight, or more generally for groups of columns associated with the same effective penalty level, an equalcorrelation set can be defined in a manner analogous to the unweighted case. As a result, the framework of Theorem 3.1 and the main analysis extend directly, and the corresponding DMSE calculations remain valid with only notational modifications.
@@ 5 Conclusions
This paper revisits the classical tuning paradigm for the Lasso through direct comparisons of prediction risk. We show that, in certain regimes, the Lasso equipped with the conventional choice $\lambda\asymp\sigma\sqrt{\log(p)/n}$ can be dominated in mean squared prediction error. Our analysis is based on an explicit DMSE decomposition, which identifies settings in which stronger regularization leads to suboptimal prediction performance. From this viewpoint, these phenomena provide a quantitative explanation for empirical observations that smaller tuning parameters may perform favorably in correlated designs and some other settings.
Although the Lasso serves as a primary reference, the arguments developed here are not specific to a single estimator and point to a broader perspective on $\ell_{1}$-regularized methods. Extensions to more general settings, including models with dependent or heteroscedastic noise and nonlinear frameworks, remain an interesting direction for future work.
@@ Appendix A Proofs of Technical Lemmas
Lemma A.2 (Newmann Serious Expansion).
Let $A\in\mathbb{R}^{d\times d}$ with $\|A\|_{\infty}<1$. Then we have
$(I-A)^{-1}=\sum_{k=0}^{\infty}A^{k}.$
For more details, see (Gentle, 2024, Section 3.11.6, p. 197).
Lemma A.3 (Closed-form Representation under the KKT Conditions).
@@
Following from the second definition of our refitting step, we have
$\frac{1}{n}X_{E}^{\top}(y-X\hat{\beta}_{\mathrm{L}})=\lambda_{\mathrm{L}}s,$
Lemma A.4 (Alternative Form of $\|\hat{\delta}_{E}\|_{1}$).
Proof.
Notice that, it suffices to show that $\operatorname{sign}(s)=\operatorname{sign}(\hat{\delta}_{E})$. According to lemma A.3, we have:
Thus, the problem can be reformulated as follows:
$\|\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}\|_{\infty}<\frac{1}{2}.$
Applying Newmann series expansion (Lemma A.2), we get:
Thus, we obtain
and our target is reduced to showing that
Noting that $s$ is a binary vector, we have
Lemma A.5 (Expectation of a Gaussian linear form under polyhedral truncation).
Recall that we have the following definition:
We have derived the following result in equation (3) of the main text:
$\Delta(\hat{\beta}_{\mathrm{L}},\hat{\beta}_{\mathrm{R}})=H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})-\frac{2}{n}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle.$
Notice that both $E_{i}$ and $s_{i}$ are determined by $\lambda_{\mathrm{L}}$, and this dependence is implicit in the notation $H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{D}})$. Moreover, on the event where $E_{i}$ and $s_{i}$ remain constant, this term is also determined as a constant. Recall the polyhedral lemma (Lee et al., 2016), let $\{(E_{i},s_{i})\}_{i=1}^{3^{p}}$ be an enumeration of all possible support–sign pairs, we have
$\mathbf{1}_{\{E=E_{i},\,s=s_{i}\}}=\mathbf{1}_{\{A_{i}\varepsilon\leq b_{i}\}}=\mathbf{1}_{\{\varepsilon\in\mathcal{P}_{i}\}},$
where $\mathcal{P}_{i}=\{\varepsilon:A_{i}\varepsilon\leq b_{i}\}$ denotes the polyhedron associated with the selection event $\{E=E_{i},s=s_{i}\}$. Collecting the above definitions and results, we obtain
$\mathrm{DMSE}$ $=\mathbb{E}H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})\sum_{i}\mathbf{1}_{\{E=E_{i},\,s=s_{i}\}}-\frac{2}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle\sum_{i}\mathbf{1}_{\{E=E_{i},\,s=s_{i}\}}$
$=\sum_{i}\mathbb{E}H_{E_{i},s_{i}}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})-\frac{2}{n}\sum_{i}\mathbb{E}\langle X_{E_{i}}\hat{\delta}_{E_{i}},\varepsilon\mathbf{1}_{\{E=E_{i},\,s=s_{i}\}}\rangle$
$=\sum_{i}\mathbb{E}H_{E_{i},s_{i}}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})-\frac{2}{n}\sum_{i}\mathbb{E}\langle X_{E_{i}}\hat{\delta}_{E_{i}},\varepsilon\mathbf{1}_{\{A_{i}\varepsilon\leq b_{i}\}}\rangle,$
where $H_{E_{i},s_{i}}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})$ denotes the value of $H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})$ on the corresponding event.
Appendix B Proofs of the Main Theorems
Proof of corollary 3.5
Consider the function
$f_{2}(x)=\frac{x(2x+1)}{(x+1)^{2}},\quad x\geq 0.$
The function $f_{2}(x)$ is monotone increasing, and it converges to $2$ as $x\to\infty$. Taking this back to Theorem 2.2 concludes the proof.
@@ Appendix B Proofs of the Main Theorems
Recall that $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$ and define, for each $j\in\{1,\dots,p\}$,
$Z_{j}:=\frac{1}{n}X_{j}^{\top}\varepsilon,$
where $X_{j}\in\mathbb{R}^{n}$ denotes the $j$-th column of $X$. For each $j$, decompose
$|Z_{j}|=\big|Z_{r(j)}+(Z_{j}-Z_{r(j)})\big|\leq|Z_{r(j)}|+|Z_{j}-Z_{r(j)}|.$
@@
$\mathbb{E}\max_{j}|Z_{j}|\leq\mathbb{E}\max_{\ell\leq K}|Z_{r_{\ell}}|+\mathbb{E}\max_{j\leq p}|Z_{j}-Z_{r(j)}|.$
First, note that each $Z_{r_{\ell}}$ is centered Gaussian with variance $\sigma^{2}/n$. By Lemma 3.7 with $p$ replaced by $K$,
$\mathbb{E}\max_{\ell\leq K}|Z_{r_{\ell}}|\leq\frac{\sigma}{\sqrt{n}}\sqrt{2\log(2K)}.$
@@
$Z_{j}-Z_{r(j)}=\frac{1}{n}(X_{j}-X_{r(j)})^{\top}\varepsilon,\quad\operatorname{Var}(Z_{j}-Z_{r(j)})=\frac{\sigma^{2}}{n^{2}}\|X_{j}-X_{r(j)}\|_{2}^{2}\leq\frac{\sigma^{2}}{n}\delta^{2}.$
Thus $\{Z_{j}-Z_{r(j)}\}_{j=1}^{p}$ are centered Gaussian variables with variance at most $\sigma^{2}\delta^{2}/n$. By Lemma 3.7,
$\mathbb{E}\max_{j\leq p}|Z_{j}-Z_{r(j)}|\leq\frac{\sigma}{\sqrt{n}}\,\delta\,\sqrt{2\log(2p)}.$
Combining the two bounds completes the proof:
$\mathbb{E}\max_{1\leq j\leq p}|Z_{j}|\leq\frac{\sigma}{\sqrt{n}}\big(\sqrt{2\log(2K)}+\delta\sqrt{2\log(2p)}\big).$
@@
where and we adopt the same notation for any index set, in particular, $\Sigma_{n,S_{0}}\coloneqq X_{S_{0}}^{\top}X_{S_{0}}/n$, and $s_{0}$ denotes the sign vector corresponding to $s$ on the event $E={S}_{0}$. Note that, by symmetry of the Gaussian distribution, we have
@@
$\|X_{E}\hat{\delta}_{E}\|_{2}^{2}$ $=\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle$
@@
Consider the function
@@
We observe that $f_{3}(x)$ attains its maximum at $f(a)$. Thus, we have
@@
using a similar approach. First, following an argument similar to that used to derive inequality (3), we obtain
@@
$\frac{1}{n}\mathbb{E}[\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle\mathbf{1}_{\{E=S_{0}\}}]$
@@
$f_{3}(x)=\frac{\sqrt{x}(2x+1)}{(x+1)^{2}},\qquad x\geq 0.$
Note that $f_{3}(x)$ increases before $x\approx 0.79$, decreases after $x\approx 0.79$, and attains its maximum value $f_{3,\max}\approx 1.78$ at $x\approx 0.79$.

### v2

@@ 3 Main Results
A proof is provided in Appendix B.
Remark 3.11 (Subspace covering and induced clustering).
Suppose that the normalized columns $\{X_{j}/\sqrt{n}\}_{j=1}^{p}$ lie in an $r$-dimensional subspace of $\mathbb{R}^{n}$. Under our column normalization assumption, these points lie on the unit sphere $S^{r-1}$ within that subspace. By a standard covering argument on this sphere (see, for example, (Vershynin, 2018, Corollary 4.2.13)), for any $\delta\in(0,1]$, there exist vectors $\{v_{\ell}\}_{\ell=1}^{K}$ in the subspace, each satisfying $\|v_{\ell}\|_{2}=\sqrt{n}$, with $K\leq(1+2/\delta)^{r}$, such that
$\|X_{j}-v_{r(j)}\|_{2}\leq\delta\sqrt{n},\qquad\forall j.$
This provides a structural condition under which the assumption of Lemma 3.10 holds with $K\lesssim(1/\delta)^{r}$.
Notably, exact low-dimensional containment of the columns is generally incompatible with Assumption 2.5. However, the same clustering conclusion continues to hold under approximate low-dimensional structure: if each column $X_{j}$ admits a decomposition into a dominant component lying in an $r$-dimensional subspace and a small residual component orthogonal to that subspace, then the dominant part can be approximated by representatives as above, and the clustering bound still follows by the triangle inequality.
This phenomenon is closely related to settings with highly correlated covariates, particularly when the design exhibits low-dimensional structure and irrelevant columns lie close to the span of relevant ones, as discussed in (Dalalyan et al., 2017). Our analysis complements empirical evidence favoring smaller tuning parameters. Next, we cite a generic chaining bound, which provides a more general framework and is closely related to the structural viewpoint developed in our analysis.
Lemma 3.12 (Generic chaining bound (Talagrand, 2021)).
Let $(T,d)$ be a metric space. Suppose that $\{X_{t}:t\in T\}$ is a centered process whose increments are sub-Gaussian with respect to $d$. Then there exists an absolute constant $C>0$ such that
$\mathbb{E}\sup_{t\in T}X_{t}\leq C\,\gamma_{2}(T,d),$
where
$\gamma_{2}(T,d)=\inf_{\{T_{k}\}}\sup_{t\in T}\sum_{k\geq 0}2^{k/2}\,d(t,T_{k}),$
the infimum being taken over all admissible sequences $\{T_{k}\}$, that is, $|T_{0}|=1$ and $|T_{k}|\leq 2^{2^{k}}$ for all $k\geq 1$. See also, for example, (Vershynin, 2018) for further details on chaining methods, including Dudley’s entropy integral inequality, which provides a closely related perspective.
Remark 3.13 (Intrinsic stochastic complexity and tuning calibration).
The preceding lemmas indicate that the structure of the design matrix influences the range of Lasso tuning parameters for which prediction suboptimality occurs. In particular, in regimes where the containment error $\mathbb{P}(E\not\subseteq T_{0})$ is small, the lower bound on the DMSE increases as the term $\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}/n]$ decreases, where $T_{0}$ denotes a high-probability containment set for the Lasso-selected support. This suggests that the main situations leading to pronounced prediction suboptimality can be summarized as follows:
  • a low noise level;
  • a sparse true support, and the probability of false inclusions by the Lasso is small, or more generally, the Lasso-selected support is contained in a relatively small index set with high probability;
  • strong column correlations, or more precisely, a design with significant geometric redundancy, where many directions are governed by a smaller intrinsic set, or, more delicately, by refined chaining-based control.
Consequently, the lower bound provides a finite-sample explanation for why the universal Lasso tuning level $\lambda\asymp\sigma\sqrt{(\log p)/n}$ may be overly conservative in regimes where the effective stochastic complexity is governed by a smaller containment set $T_{0}$ and fewer representative directions. Finally, we consider the case where the Lasso recovers the true support with high probability under symmetric noise and establish a complementary theorem.
@@
Let $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$, and let $S_{0}$ denote the true support. Consider our Lasso-Ridge method. Select the tuning parameter of the second step as $\lambda_{\mathrm{R}}=c|E|$, then we have
@@
This theorem shows that when the Lasso selects the true support with high probability, its prediction performance can be improved by a suitable refinement over a wide range of tuning parameters. Moreover, it is known that variable selection consistency may occur when the Lasso tuning parameter does not vanish; see, for example, (Lahiri, 2021). Besides, under standard sign-consistency conditions such as the irrepresentable condition together with a beta-min assumption, the model selection error probability $\mathbb{P}(E\neq S_{0})$ becomes exponentially small; see (Zhao and Yu, 2006). In such settings, the present theorem implies that prediction improvement can be achieved. For completeness, the proof is given in Appendix B.
Remark 3.16 (Support stability beyond exact recovery).
The theorem does not require exact support recovery, i.e., $\mathbb{P}(E=S_{0})\to 1$. It suffices that the Lasso-selected support coincides with a fixed index set $S^{\star}$ with high probability, for instance, the support obtained in the noiseless case. Consequently, the refinement step can exploit this support stability to yield prediction improvement.
4 Discussion and Possible Extensions
Extensions to $\ell_{1}$-based methods.
The analysis developed in this paper is not specific to the classical Lasso, but depends more broadly on structural features shared by $\ell_{1}$-based procedures. At a high level, the extension follows the same general template: identify an intrinsic active or equicorrelation-type set induced by the regularization, perform the refitting step on this set, and derive a prediction dominance inequality from the definition of the second step, after which the main DMSE framework can be applied on the resulting intrinsic set.
This perspective is expected to extend, at least formally, to other $\ell_{1}$-based methods such as the Dantzig selector (Candès and Tao, 2007) and the Adaptive Lasso (Zou, 2006). For the Dantzig selector, one may analogously introduce an intrinsic active set associated with the constraint structure, while for the Adaptive Lasso with fixed weights, the same geometric decomposition can be adapted to the weighted $\ell_{1}$ setting, leading to a corresponding equicorrelation-type structure. A complete extension to these methods, especially for a precise DMSE analysis, requires additional technical developments and is therefore left for future work. Next, we discuss extensions to martingale noise structures.
Martingale noise.
The results of this paper do not fundamentally rely on independence or Gaussianity of the noise. In particular, the framework also accommodates martingale difference noise under suitable moment or conditional variance assumptions. Consider the fixed-design linear regression model $y_{i}=x_{i}^{\top}\beta_{0}+\varepsilon_{i}$, $i=1,\dots,n$, where $X\in\mathbb{R}^{n\times p}$ is deterministic with normalized columns $\|X_{j}\|_{2}^{2}=n$ for all $j=1,\dots,p$, and $(\varepsilon_{i},\mathcal{F}_{i})_{i=1}^{n}$ is a martingale difference sequence satisfying $\mathbb{E}(\varepsilon_{i}\mid\mathcal{F}_{i-1})=0$ and $\mathbb{E}(\varepsilon_{i}^{2}\mid\mathcal{F}_{i-1})\leq\sigma^{2}$ almost surely. Assume further that the design obeys a mild regularity condition such as $\max_{i,j}|x_{ij}|\leq K$. Then, for any index set $T$ and each $j\in T$,
$(X_{T}^{\top}\varepsilon)_{j}=\sum_{i=1}^{n}x_{ij}\varepsilon_{i}$
is a weighted sum of martingale differences. Hence, by Theorem A1.1.6 in (Nishiyama, 2021, p. 219), $\mathbb{E}\|X_{T}^{\top}\varepsilon/n\|_{\infty}$ is of order $\sqrt{\log|T|/n}$, up to constants. Substituting such a maximal inequality into Theorem 3.1 yields a corresponding range of tuning parameters for which the Lasso may exhibit suboptimal prediction performance under martingale noise. Such settings arise naturally in heteroskedastic regression, econometrics, and time-series models with predictable conditional variances.
@@ 5 Conclusions
This paper revisits the tuning paradigm for the Lasso from the perspective of prediction risk dominance. In particular, we propose tuning ranges in which suboptimality arises; more specifically, the resulting predictor can be dominated in mean squared prediction error by refitting on the Lasso-selected support. From this perspective, the results provide a quantitative explanation for empirical observations that smaller tuning parameters may perform favorably in correlated designs and related settings, and that refitting on the selected support of the Lasso may improve prediction performance.
Although the Lasso serves as the primary reference point, the arguments rely mainly on the structure of $\ell_{1}$-regularization and the associated DMSE decomposition, rather than on features unique to a specific estimator. Extending the analysis to more general settings, such as nonlinear models, is left for future work. Developing new frameworks for the standard linear regression model is also an interesting direction.
@@ Appendix A Proofs of Technical Lemmas
Lemma A.1 (Neumann Serious Expansion).
Let $A\in\mathbb{R}^{d\times d}$ with $\|A\|_{\infty}<1$. Then we have
$(I-A)^{-1}=\sum_{k=0}^{\infty}A^{k}.$
For more details, see, for example, (Gentle, 2024, Section 3.11.6, p. 197).
Lemma A.2 (Closed-form Representation under the KKT Conditions).
Proof.
Following from the second definition of our refitting step, we have
$\frac{1}{n}X_{E}^{\top}(y-X\hat{\beta}_{\mathrm{L}})=\lambda_{\mathrm{L}}s.$
Lemma A.3 (Alternative Form of $\|\hat{\delta}_{E}\|_{1}$).
Proof.
Notice that, it suffices to show that $\operatorname{sign}(s)=\operatorname{sign}(\hat{\delta}_{E})$. According to lemma A.2, we have
Thus, the problem can be reformulated as follows:
$\|\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}\|_{\infty}<\frac{1}{2}.$
Applying Neumann series expansion (Lemma A.1), we get
Thus, we obtain
and our target is reduced to showing that
Noting that $s$ is a binary vector, we have
@@
Lemma A.5.
Assume that $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$. Then
$\mathbb{P}(E=\emptyset)\leq\Phi\!\left(\frac{\lambda_{\mathrm{L}}-\|\Sigma_{n}\beta_{0}\|_{\infty}}{\sigma/\sqrt{n}}\right),$
where $\Sigma_{n}:=X^{\top}X/n$.
@@
By the KKT conditions, $\hat{\beta}_{\mathrm{L}}=0$ if and only if
$\frac{1}{n}\,\|X^{\top}y\|_{\infty}\leq\lambda_{\mathrm{L}}.$
Hence
$\mathbb{P}(E=\emptyset)$ $=\mathbb{P}(\hat{\beta}_{\mathrm{L}}=0)$
$=\mathbb{P}\!\left(\frac{1}{n}\,\|X^{\top}y\|_{\infty}\leq\lambda_{\mathrm{L}}\right).$
Let $j^{\star}$ satisfy
$|(\Sigma_{n}\beta_{0})_{j^{\star}}|=\|\Sigma_{n}\beta_{0}\|_{\infty}.$
Then
$\left\{\frac{1}{n}\,\|X^{\top}y\|_{\infty}\leq\lambda_{\mathrm{L}}\right\}\subseteq\left\{\frac{1}{n}\,|X_{j^{\star}}^{\top}y|\leq\lambda_{\mathrm{L}}\right\}.$
Write
$m:=(\Sigma_{n}\beta_{0})_{j^{\star}}=\frac{1}{n}X_{j^{\star}}^{\top}X\beta_{0},\qquad|m|=\|\Sigma_{n}\beta_{0}\|_{\infty}.$
Using $y=X\beta_{0}+\varepsilon$, we obtain
$\mathbb{P}(E=\emptyset)\leq\mathbb{P}\!\left(\left|\frac{1}{n}X_{j^{\star}}^{\top}\varepsilon+m\right|\leq\lambda_{\mathrm{L}}\right).$
Let
$Z:=\frac{1}{n}X_{j^{\star}}^{\top}\varepsilon.$
Since $Z\sim N(0,\sigma^{2}/n)$ is symmetric about zero, we have $Z\stackrel{{\scriptstyle d}}{{=}}-Z$. Consequently,
$\mathbb{P}(|Z+m|\leq\lambda_{\mathrm{L}})$ $=\mathbb{P}(|-Z+m|\leq\lambda_{\mathrm{L}})$
$=\mathbb{P}(|Z-m|\leq\lambda_{\mathrm{L}}).$
If $m\geq 0$, then |Zm|=|Z|m|||Z-m|=|Z-|m||; if $m<0$, then m=|m|-m=|m| and |Zm|=|Z+|m|||Z-m|=|Z+|m||. In either case,
$\mathbb{P}(|Z+m|\leq\lambda_{\mathrm{L}})=\mathbb{P}(|Z+|m||\leq\lambda_{\mathrm{L}}).$
Because $|m|\geq 0$, the event $\{|Z+|m||\leq\lambda_{\mathrm{L}}\}$ is contained in $\{Z\leq\lambda_{\mathrm{L}}-|m|\}$, and therefore
$\mathbb{P}(E=\emptyset)\leq\mathbb{P}\!\left(Z\leq\lambda_{\mathrm{L}}-|m|\right)=\Phi\!\left(\frac{\lambda_{\mathrm{L}}-\|\Sigma_{n}\beta_{0}\|_{\infty}}{\sigma/\sqrt{n}}\right).$
Remark A.6.
In particular, if $\|\Sigma_{n}\beta_{0}\|_{\infty}\geq\lambda_{\mathrm{L}}+t\,\sigma/\sqrt{n}$ for some $t>0$, then
$\mathbb{P}(E=\emptyset)\leq\Phi(-t),$
which decays exponentially in $t^{2}$. Hence, under the usual scaling in which $\lambda_{\mathrm{L}}\to 0$ while $\|\Sigma_{n}\beta_{0}\|_{\infty}$ remains of constant order, the margin $\|\Sigma_{n}\beta_{0}\|_{\infty}-\lambda_{\mathrm{L}}$ is bounded away from zero, and consequently the probability that the Lasso solution is exactly zero is exponentially small.
Appendix B Proofs of Auxiliary Results
@@
Recall that $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$ and, for each $j\in\{1,\dots,p\}$,
$Z_{j}=\frac{1}{n}X_{j}^{\top}\varepsilon.$
For each $j$, decompose
$|Z_{j}|=|\frac{1}{n}v_{r(j)}^{\top}\varepsilon+(Z_{j}-\frac{1}{n}v_{r(j)}^{\top}\varepsilon)|\leq|\frac{1}{n}v_{r(j)}^{\top}\varepsilon|+|Z_{j}-\frac{1}{n}v_{r(j)}^{\top}\varepsilon|.$
@@
$\mathbb{E}\max_{j}|Z_{j}|\leq\mathbb{E}\max_{\ell\leq K}|\frac{1}{n}v_{\ell}^{\top}\varepsilon|+\mathbb{E}\max_{j\leq p}|Z_{j}-\frac{1}{n}v_{r(j)}^{\top}\varepsilon|.$
First, each $v_{\ell}^{\top}\varepsilon/n$ is centered Gaussian with variance
$\frac{\sigma^{2}}{n^{2}}\|v_{\ell}\|_{2}^{2}\leq\frac{\sigma^{2}}{n}.$
Applying Lemma 3.8 with $p$ replaced by $K$, we obtain
$\mathbb{E}\max_{\ell\leq K}|\frac{1}{n}v_{\ell}^{\top}\varepsilon|\leq\frac{\sigma}{\sqrt{n}}\sqrt{2\log(2K)}.$
@@
$Z_{j}-\frac{1}{n}v_{r(j)}^{\top}\varepsilon=\frac{1}{n}(X_{j}-v_{r(j)})^{\top}\varepsilon,\quad\operatorname{Var}(Z_{j}-\frac{1}{n}v_{r(j)}^{\top}\varepsilon)\leq\frac{\sigma^{2}}{n}\delta^{2}.$
Thus $\{Z_{j}-v_{r(j)}^{\top}\varepsilon/n\}_{j=1}^{p}$ are centered Gaussian variables with variance at most $\sigma^{2}\delta^{2}/n$. By Lemma 3.8,
$\mathbb{E}\max_{j\leq p}|Z_{j}-\frac{1}{n}v_{r(j)}^{\top}\varepsilon|\leq\frac{\sigma}{\sqrt{n}}\,\delta\,\sqrt{2\log(2p)}.$
Combining the two bounds completes the proof. ∎
@@
where we adopt the same notation for any index set. In particular, $\Sigma_{n,S_{0}}\coloneqq X_{S_{0}}^{\top}X_{S_{0}}/n$, and $s_{0}\coloneqq\operatorname{sign}(\beta_{0,S_{0}})$ denotes the (deterministic) sign vector on the event $E=S_{0}$. Note that the vector $\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0}$ is non-random; hence, by symmetry of the Gaussian noise $\varepsilon$, we have
@@
$\frac{1}{n}\|X_{E}\hat{\delta}_{E}\|_{2}^{2}$ $=\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle$
@@
$=\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{\mu_{1}}{(\mu_{1}+\lambda_{\mathrm{R}})^{2}},\dots,\frac{\mu_{|E|}}{(\mu_{|E|}+\lambda_{\mathrm{R}})^{2}})P^{\top}s,$
where $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{|E|}\geq 0$ denote the ordered eigenvalues of $\Sigma_{n,E}=X_{E}^{\top}X_{E}/n$ and $P$ is the corresponding eigenvector matrix. Consider the function
@@
Note that $f_{3}(x)$ attains its maximum at $x=a$. Thus, we have
@@
using a similar approach. First, note that on the event $E=S_{0}$, we have
$\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle=\frac{1}{n}\|\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0}\|_{2}^{2}.$
Following an argument similar to that used to derive inequality (3), we obtain
@@
$f_{4}(x)=\frac{\sqrt{x}(2x+1)}{(x+1)^{2}},\qquad x\geq 0.$
A direct calculation shows that $f_{4}(x)$ increases on $0,(3+\sqrt{17})/4]$, decreases thereafter, and attains its maximum value $f_{4,\max}\approx 0.79$ at $x\approx 1.78$.
Acknowledgement
The author gratefully acknowledges Professor Yoichi Nishiyama for his valuable guidance, insightful comments, and technical assistance.